Solve the following quadratic equations:

1) 2x2+ 6x+ 1 = 0

2) 4x2+ 12x+ 2 = 0

3) 9x2−4x+ 12 = 0

4) (x−1)(x+ 3 +3(x−1))=(4−x)(x−1)(x−1)

Hint:Factor and simplify using (x−1).
Answer #1

Gtven Equating it to stamdlar dl dorm af uaoleahc ez ax2tb04co -btvb-4ac aa 6t36-4 2 1 a. 2 628 -6 36- 8 Z-6t5. 29 24 0.l77Gr47212 = O uith stanaleTd guadra c ezuah en Equaing a-9, b-4, C-12 We tenouo +hat - bt Vbuac 2 a 4t 16-4.g.12 2.9 4t J-416 4a) Ca-1) Cz42 + 3 Cx-)) = C-x) C2-)(2-1) Xt 3 D(x-1) 3x-3 Cx-1uR 4-2) (217 teRe Cu-) Cemmon C-1) Ce-l)un - Inz521 Cn-i 22-2ty